Problem 3: Public Goods sumers, Ben and Joe. There is public good in this economy in a P 10- Q, and Joes demand sirens in the markets is form of tornado sirens. Bens demand for tornado sirens is given by raso sirens in the markets is for tornado sirens is P 8 2Q. Marginal cost for providing torna constant, MC 9. a) Are tornado sirens non-exclusive? Explain your answer. Is there a potential for free-rider problem? Derive market demand curve for sirens. Draw three graphs on the top of each other- first graph for Bens demand, second graph for Joes demand, and the third graph for b) C) market demand. d) How many sirens will be provided in the market? e) What will be the price for these sirens? f is the result from previous question realistic? Discuss how government funds public goods.
Answer #1

a) Public goods are the goods which are non-rival and non-exclusive in nature. Non-rival means consumption of the public good by one person does not reduce its amount available for the other person. For example National defense, public park, street light etc. Non-exclusion means if public good has been provided then it does not possible to exclude one person from consuming it.

Tornado sirens are non-exclusive.

b) Yes.

Public good is a good which are non-rival and non-exclusive in nature. Non-rival means consumption of the public good by one person does not reduce its amount available for the other person. Non-exclusion means if public good has been provided then it does not possible to exclude one person from consuming it. In fact it will be Pareto inefficient to exclude one person because if that person is allowed to consume then he may be better off but other are not worse off. As a result of non-rival and non-exclusion, property there is problem of free rider and due to free rider market fails.

c) Ben's demand; P = 10 - Q

Q = 10 - P

Joe's demand; P = 8 - 2Q

Q = (8 - P)/2 = 4 - 0.5P

Market demand = Ben's demand + Joe's demand = 10 - P + 4 - 0.5P

Qd = 14 - 1.5P

oe oO P O9.33 lo 6 ч81216 Maoket o y812 mc - EquimMR mc

CK e) 9.17 e)